Find the equation of a line thru (-3,-12) and parallel to the line thru(6,0) and(4,-5).
Answer by rfer(16322)

m=-5-0/4-6
m=-5/-2
m=2.5
---------------
y=2.5x+b
-12=2.5*-3+b
-12=-7.5+b
b=-4.5
-----------------
y=2.5x-4.5